### v2

# RefineCoder: Iterative Improving Large Language Models through Adaptive Critique Refinement for Code Generation

Abstract
Code generation has attracted increasing attention with the rise of Large Language Models (LLMs). Many studies have developed powerful code LLMs by synthesizing code-related instruction data and applying Supervised Fine-Tuning (SFT). However, these methods are limited by teacher model distillation and ignore the potential of iterative refinement by self-generated code. In this paper, we propose Adaptive Critique Refinement (ACR), which enables the model to refine itself by self-generated code response and teacher-generated critique, rather than directly imitating the teacher-generated code responses. Concretely, ACR includes a composite scoring system with LLM-as-a-Judge to evaluate the quality of code responses and a selective critiquing strategy with LLM-as-a-Critic to critique self-generated low-quality code responses. We develop the RefineCoder series by iteratively applying ACR, achieving continuous performance improvement on multiple code generation benchmarks. Compared to the baselines of the same size, our proposed RefineCoder series can achieve superior performance using less data.
RefineCoder: Iterative Improving Large Language Models through Adaptive Critique Refinement for Code Generation
Changzhi Zhou1thanks:   Work done during internship at Meituan. Equal contribution.,   Xinyu Zhang [^1] footnotemark: 1,   Dandan Song1thanks:   Corresponding Author.,  Xiancai Chen2, Wanli Gu3, Huipeng Ma1, Yuhang Tian1, Mengdi Zhang3, Linmei Hu122footnotemark: 2 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China 2Peking University   3Meituan zhouchangzhi97@gmail.com   {sdd,hulinmei}@bit.edu.cn
1 Introduction
Code generation, also called program synthesis, is a key application area of Large Language Models (LLMs) and has attracted significant attention from the research community Gulwani et al. (2017); Chen et al. (2021). Numerous studies focus on developing code-specific pre-trained models, such as CodeLlama Roziere et al. (2023), DeepSeek-Coder Guo et al. (2024), and Qwen-Coder Hui et al. (2024). Despite their strong capabilities in code understanding, these LLMs typically require additional post-training to become effective and user-friendly for code-related tasks.
Figure 1: Comparison of two fine-tuning paradigms. DFT uses code distilled by teacher model, whereas ACR adaptively constructs distinct data formats infusing self-generated code and teacher-generated critique.
As a popular post-training technology, Distillation-based Fine-Tuning (DFT) leverages code instruction data synthesized by powerful teacher models (often proprietary LLMs like GPT-4) to fine-tune code LLMs, as shown in Figure 1. Many works Zhang et al. (2024b) based on this paradigm have been proposed recently. For example, Code Alpaca Chaudhary (2023) first establishes this synthesis paradigm using Self-Instruct Wang et al. (2023). Subsequently, Code-related Evol-Instruct Luo et al. (2024) prompts LLMs to evolve more intricate code instructions, while OSS-Instruct Wei et al. (2024b) generates more realistic code instructions based on open-source code snippets Kocetkov et al. (2022). In contrast, OpenCodeInterpreter Zheng et al. (2024a) enables code LLMs to learn from compiler diagnostics and human feedback by constructing multi-turn feedback data. These diverse synthesis strategies have significantly advanced the development of code LLMs. However, this paradigm of teacher model distillation inevitably suffers from a critical limitation: the performance of the student model largely relies on the teacher model, ignoring the potential of iterative refinement by self-generated code.
Apart from research on code instruction fine-tuning, some other studies Chen et al. (2023, 2024, 2025); Jiang et al. (2025) have explored the ability of LLMs to refine self-generated code during inference. For example, Self-Debugging Chen et al. (2024) iteratively debugs self-generated code using feedback from the code executor. However, these methods that freeze parameters are essentially a form of prompt engineering with external feedback. They cannot improve the intrinsic one-off code generation ability of LLMs. Besides, multiple calls to LLMs increase inference latency.
In this paper, we propose a novel fine-tuning paradigm called Adaptive Critique Refinement (ACR). It evolves through a combination of self-generated code and teacher-driven critique, thereby enhancing its capacity for one-shot code generation and surpassing the limitations of conventional teacher distillation methods. Specifically, as shown in Figure 1, ACR evaluates both self-generated and teacher-generated code, adaptively critiques low-quality self-generated responses, and generates new samples based on the evaluation and critique results. This process is analogous to a student initially attempting to solve problems independently and later comparing their solutions against given answers, rather than simply copying the provided solutions. By iteratively applying the ACR process, the model achieves continuous improvement in its code generation capabilities. Notably, we design a composite scoring system and a selective critique strategy to perform scoring and critique, respectively. The scoring system combines LLM-as-a-Judge and a code executor to provide accurate and comprehensive evaluations. The critique strategy adaptively constructs two types of data and incorporates LLM-as-a-Critic.
We conduct iterative ACR based on DS-Coder-6.7B-Base Guo et al. (2024) and Qwen2.5-Coder-7B-Base Hui et al. (2024), leading to the development of the RefineCoder series. After three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B demonstrate significant improvements in average pass@1 performance on the HumanEval(+), MBPP(+), LiveCodeBench, and BigCodeBench (hard) Chen et al. (2021); Austin et al. (2021); Liu et al. (2023); Jain et al. (2024); Zhuo et al. (2024). The key contributions of this paper are as follows:
1) We propose Adaptive Critique Refinement (ACR), a novel fine-tuning paradigm that refines code LLMs using self-generated code and external critique, on the basis of which we develop a series of strong code LLMs namely RefineCoder.
2) To ensure the effectiveness of ACR, we design a composite scoring system with LLM-as-a-Judge and a selective critiquing strategy with LLM-as-a-Critic, enabling both accurate scoring and critique of code responses.
3) Experimental results from the RefineCoder series show that iterative ACR continuously improves code generation performance. After three iterations, the RefineCoder models outperform baselines of the same size while requiring less data.
2 Related Work
LLMs for Code Generation LLMs have shown exceptional code understanding abilities due to extensive pre-training on code-related corpora. Numerous models like GPT-4o OpenAI (2024a), Gemini Google (2024), Claude-3.5 Anthropic (2024), Qwen2.5-Coder Hui et al. (2024), and DeepSeek-Coder Guo et al. (2024) exhibit strong performance on code generation benchmarks. Recently, the release of OpenAI o1 OpenAI (2024b) and DeepSeek-R1 DeepSeek-AI (2025) has spurred a surge in research on deep reasoning LLMs, achieving expert-level performance on competitive programming problems (e.g., CodeForces) and further advancing LLMs in the code domain.
Figure 2: Overview of ACR in the $t$-th iteration. (1) Sampling: The model $M_{t}$ samples $n$ code responses $\{y^{i}\}_{i=1}^{n}$ with high temperature for a programming instruction $x$ in the dataset $D_{t}$. (2) Ranking: A composite scoring system first scores all self-generated code responses $\{y^{i}\}_{i=1}^{n}$, then selects the highest-scoring response for pairwise comparison with the original code response $y^{0}$ to identify the winner $y^{w}$ and the loser $y^{l_{1}}$. (3) Refining: Depending on the identity of the $y^{w}$, the selective critiquing strategy constructs a new single-turn data $(x,y^{w})$ or two-turn critique data $(x,y^{l_{1}},c,y^{w})$, where $c$ is critique. The original data $(x,y)$ is replaced with new data and all the new data form a new dataset $D_{t+1}$. (4) Training: The new model $M_{t+1}$ is fine-tuned using $D_{t+1}$.
Distillation-based Code Fine-Tuning Unlike proprietary models, many studies focus on fine-tuning open-source code pre-trained models, which has greatly contributed to the rapid development of the code generation field. A key technique for achieving this is distilling data from teacher models. Code Alpaca Chaudhary (2023) introduces Self-Instruct Wang et al. (2023) to distill GPT-3. Following this, WizardCoder Luo et al. (2024) evolves more complex code instruction data using Evol-Instruct Xu et al. (2024). MagiCoder Wei et al. (2024b) proposes OSS-Instruct, where the model generates instructions and code responses sequentially based on open-source code snippets, thus creating more practical code data. In contrast to OSS-Instruct, InverseCoder Wu et al. (2024) reverses the order of instruction and code response generation. WaveCoder Yu et al. (2024) constructs a multi-task code dataset, enhancing the model’s versatility. Besides, OpenCodeInterpreter Zheng et al. (2024a) builds multi-turn code data, enabling the model to learn from execution feedback and human feedback. However, these methods aim to enable the model to learn by imitating the teacher, overlooking the potential for refinement through self-generated code.
Iterative Self-Refinement of LLMs Iterative self-refinement refers to LLMs enhancing themselves iteratively by utilizing self-generated responses with the help of external signals. One line of research Huang et al. (2023); Madaan et al. (2023); Hu et al. (2024) focuses on enabling self-correction during the inference stage by iteratively calling LLMs and incorporating external signals. In the code domain, CodeT Chen et al. (2023), Self-Debugging Chen et al. (2024), and LDB Zhong et al. (2024) follow this approach. However, these prompt engineering methods with external feedback cannot improve the intrinsic capabilities. Another line of research focuses on iteratively training the model using self-generated outputs to enhance its intrinsic capabilities Dong et al. (2025); Yuan et al. (2024); Yan et al. (2025); Kim et al. (2025). These methods typically rely on preference learning, such as DPO Rafailov et al. (2023). CodeLutra Tao et al. (2024) has successfully applied this approach to the code domain, but it heavily depends on golden labels, which limits its applicability. In contrast to the works above, we propose the ACR method, which achieves iterative self-refinement by only using a simple Supervised Fine-Tuning (SFT) loss. Our approach is orthogonal to prompt engineering methods like Self-Debugging and is more efficient and generalizable than CodeLutra.
3 Methodology
3.1 Overview
Adaptive Critique Refinement (ACR) aims to improve code LLMs by refining the existing dataset with self-generated code and external critique. The iterative application of ACR facilitates continuous improvement in code generation performance. As illustrated in Figure 2, ACR starts with a current dataset and a code model fine-tuned on it. The method updates this dataset through a process of sampling, ranking, and refining. The updated dataset is used for the next round of fine-tuning, iteratively improving the model’s code generation capabilities.
In the following sections, we will introduce two key components of the pipeline above: the composite scoring system for ranking and the selective critiquing strategy for data refining. Finally we describe the iterative training and the resulting RefineCoder model.
3.2 Composite Scoring System with LLM-as-a-Judge
A comprehensive and accurate evaluation of code response quality is the foundation for effective ACR. Previous works Chen et al. (2023); Zhang et al. (2024a) rely on generating test cases to evaluate code quality. However, this method has issues with inaccurate test case generation Jain et al. (2024), inability to handle special judge Quan et al. (2025), and failure to evaluate non-code content in response such as code comments and explanations.
Drawing on the judgment and evaluation abilities of LLMs Dong et al. (2025); Li et al. (2024), we employ LLM-as-a-Judge as the backbone of the scoring system, supplemented by a code executor to ensure effective evaluation. Specifically, we perform a multi-faceted, point-wise judgment of each self-generated code response $y^{i}$ ($1\leq i\leq n$):
$s^{i}_{1},s^{i}_{2},...,s^{i}_{m}=\textrm{Multi-Faceted Judge}(x,y^{i}),$ (1)
where $s^{i}_{j}$ denotes the score for $y^{i}$ under the $j$-th evaluation criterion, and $m$ is the total number of criteria (e.g., code correctness, clarity). Next, we calculate the final point-wise score for each self-generated code response by computing the weighted sum and normalizing the result:
$s^{i}$ $=\sum_{j=1}^{m}\alpha_{j}s_{j}^{i},$ (2)
$\text{score}^{i}_{\text{point}}$ $=\frac{s^{i}-s_{\min}}{s_{\max}-s_{\min}},$ (3)
where $\alpha_{j}$ is the weight of the $j$-th criterion, and $s_{\min}$ and $s_{\max}$ are the minimum and maximum scores across self-generated code responses. In addition, we use a code executor to evaluate the executability of all code responses (including $y^{0}$):
$\textrm{score}_{\text{exec}}^{i}=\textrm{Executor}(y^{i}),$ (4)
where $i\in[0,n]$ and $\textrm{score}_{\text{exec}}^{i}\in\{0,1\}$. The highest-scoring self-generated code is obtained by summing the point-wise score and execution score:
$y^{self}=\arg\max_{i}(\text{score}^{i}_{\text{point}}+\text{score}^{i}_{\text{exec}}).$ (5)
where $i\in[1,n]$. Finally, we determine the winner between the best self-generated code $y^{self}$ and the teacher-generated code $y^{0}$ based on the pairwise judgment and execution scores:
$\textrm{score}_{\text{pair}}^{self},\textrm{score}_{\text{pair}}^{0}=\textrm{PairJudge}(x,y^{self},y^{0}),$ (6)
$\textrm{score}^{i}=\textrm{score}^{i}_{\text{pair}}+\textrm{score}^{i}_{\text{exec}},$ (7)
$y^{w},y^{l_{1}}=\begin{cases}y^{self},y^{0},&\text{if }\textrm{score}^{self}>\textrm{score}^{0},\\ y^{0},y^{self},&\text{if }\textrm{score}^{0}>\textrm{score}^{self},\end{cases}$ (8)
where $i\in\{self,0\}$, $\textrm{score}^{i}_{pair}\in[0,1]$, and $y^{w}$ and $y^{l_{1}}$ denote the winner and loser, respectively. The prompts for point-wise and pair-wise judgment are shown in Figures 8 and 9.
3.3 Selective Critiquing Strategy with LLM-as-a-Critic
After selecting winner and loser based on quality, we design the selective critiquing strategy as the data construction engine of ACR. It effectively utilizes high-quality and low-quality self-generated code in distinct ways, while further providing external critique for the low-quality data. This strategy mirrors the process of a student first solving problems and then critiquing their answers based on provided solutions rather than directly imitating the solutions Wang et al. (2025).
Concretely, when all self-generated code responses are of lower quality than the original code in the current dataset (i.e., $y^{w}$ is $y^{0}$ from origin dataset and $y^{l_{1}}$ is the best self-generated code response), LLM-as-a-Critic is used to critique $y^{l_{1}}$, using $y^{0}$ as a reference (The prompt is shown in Figure 10):
$c=\textrm{Critic}(x,y^{w},y^{l_{1}}),$ (9)
where $c$ is the critique of $y^{l_{1}}$. Otherwise, if $y^{w}$ is the self-generated code response, we directly construct new single-turn data. The data update rules are as follows:
$(x,y)\longrightarrow\begin{cases}(x,y^{l_{1}},c,y^{w}),&\text{if }y^{w}\text{ is }y^{0},\\ (x,y^{w}),&\text{otherwise}.\end{cases}$ (10)
If the data to be updated comes from a two-turn data rather than a single-turn one, only the instruction and the final code response will be considered. The selective critiquing strategy updates all data to comprise a new instruction dataset $D_{t+1}$.
3.4 Iterative Training
The previous sections describe the process of a single ACR in Figure 2. Next, we define the RefineCoder in the iterative ACR process:
$\circ$ RefineCoder Iter0 ($M_{0}$): The code model is trained using the initial SFT dataset $D_{0}$.
$\circ$ RefineCoder Iter1 ($M_{1}$): The code model is trained using dataset $D_{1}$, where $D_{1}$ is obtained by refining $D_{0}$ using $M_{0}$.
$\circ$ RefineCoder Itert ($M_{t}$): The code model is trained using dataset $D_{t}$, where $D_{t}$ is obtained by refining $D_{t-1}$ using $M_{t-1}$.
We fine-tune the original pre-trained model from scratch in each iteration to prevent overfitting, consistent with previous iterative training works Wang et al. (2024a); Dong et al. (2025).
Models Data Size LiveCodeBench BigCodeBench-hard Average
v3-v5 v5-v6 Instruct Complete
[Uncaptioned image] DSCoder-6.7B-Base / / / / 13.5 /
DSCoder-6.7B-Instruct 2BT 12.8 14.0 10.1 15.5 13.1
MagiCoder-S-DS-6.7B 185K 13.4 14.6 13.5 12.8 13.6
OpenCodeInterpreter-DS-6.7B 178K 7.3 15.2 13.5 16.9 13.2
WaveCoder-Ultra-6.7B 130K 12.9 14.5 12.8 16.9 14.3
RefineCoder-DS-6.7B   Iter0 80K 13.1 15.5 18.2 10.1 14.2
RefineCoder-DS-6.7B   Iter1 80K 14.3 17.2 18.9 14.2 16.2↑2.0
RefineCoder-DS-6.7B   Iter2 80K 14.3 17.8 19.6 14.2 16.5↑2.3
RefineCoder-DS-6.7B   Iter3 80K 14.2 17.9 20.1 15.5 16.9↑2.7
[Uncaptioned image]Qwen2.5-Coder-7B-Base / / / / 16.2    /
Qwen2.5-Coder-7B-Instruct Millions 18.1 18.9 20.3 20.3 19.4
RefineCoder-QW-7B   Iter0 80K 18.0 18.7 18.5 19.3 18.6
RefineCoder-QW-7B   Iter1 80K 19.6 21.3 20.7 21.8 20.9↑2.3
RefineCoder-QW-7B   Iter2 80K 20.5 21.5 20.6 22.9 21.4↑2.8
RefineCoder-QW-7B   Iter3 80K 21.7 21.9 20.8 22.3 21.7↑3.1
Table 1: Results on LiveCodeBench and BigCodeBench-hard. v3-v5 and v5-v6 refer to the questions from the periods 202407-202502 and 202502-202505, respectively. [Uncaptioned image] and [Uncaptioned image] denotes the two base models, and the rest are the instruction models. 2BT denotes 2B tokens.
4 Experiments
4.1 Benchmarks
We use HumanEval(+) Chen et al. (2021), MBPP(+) Austin et al. (2021); Liu et al. (2023) and BigCodeBench Zhuo et al. (2024) to assess the fundamental coding ability of code model. Meanwhile, we use LiveCodeBench Jain et al. (2024) and BigCodeBench-hard to measure the advanced coding ability of code model.
4.2 Baselines
$\circ$ Based on DSCoder-6.7B-Base: DSCoder-6.7B-Instruct Guo et al. (2024), MagiCoder-S-DS-6.7B Wei et al. (2024b), OpenCodeInterpreter-DS-6.7B Zheng et al. (2024a), WaveCoder-Ultra-6.7B Yu et al. (2024).
$\circ$ Based on Qwen2.5-Coder-7B-Base: Qwen2.5-Coder-7B-Instruct Hui et al. (2024).
4.3 Initial SFT Dataset $D_{0}$
The existing datasets vary in quality and carry the risk of data leakage Wang et al. (2024b); Yu et al. (2024). To ensure the accuracy and reliability of the experimental results, we construct a high-quality SFT dataset by combining Evol-Instruct and SelfCodeAlign Wei et al. (2024a). Concretely, we first prompt GPT-4o to generate code-related concepts and corresponding programming instructions from Python code snippets. These code snippets have been preprocessed to ensure high quality and contamination-free. Then, we prompt GPT-4o again to iteratively evolve these instructions using the Evol-Instruct strategy. Finally, we prompt GPT-4o to generate code responses for the instructions. Through this pipeline, we obtain an 80K high-quality and diverse python instruction dataset. See Appendix B for detailed data construction prompts.
Models Data Size HumanEval MBPP BigCodeBench Average
Base Plus Base Plus Inst Comp
[Uncaptioned image] DSCoder-6.7B-Base / 47.6 39.6 72.0 58.7 / 41.8 /
DSCoder-6.7B-Instruct 2BT 74.4 71.3 74.9 65.6 35.5 43.8 60.9
MagiCoder-S-DS-6.7B 185K 76.8 71.3 79.4 69.0 36.2 47.6 63.4
OpenCodeInterpreter-DS-6.7B 178K 77.4 72.0 76.5 66.4 37.1 44.6 62.3
WaveCoder-Ultra-6.7B 130K 75.0 69.5 74.9 63.5 33.9 43.7 60.1
RefineCoder-DS-6.7B   Iter0 80K 73.8 67.7 77.5 65.1 37.0 46.6 61.3
RefineCoder-DS-6.7B   Iter1 80K 74.4 68.9 77.0 66.4 40.6 48.0 62.6↑1.3
RefineCoder-DS-6.7B   Iter2 80K 74.4 70.3 79.6 66.9 40.6 47.9 63.3↑2.0
RefineCoder-DS-6.7B   Iter3 80K 75.0 70.7 80.2 67.2 41.1 49.6 64.0↑2.7
[Uncaptioned image]Qwen2.5-Coder-7B-Base / 61.6 53.0 76.9 62.9    /    /    /
Qwen2.5-Coder-7B-Instruct Millions 88.4 84.1 83.5 71.7 40.4 48.8 69.5
RefineCoder-QW-7B   Iter0 80K 84.9 78.1 79.4 65.8 40.0 47.4 65.9
RefineCoder-QW-7B   Iter1 80K 86.0 79.3 84.1 70.8 41.6 49.0 68.5↑2.6
RefineCoder-QW-7B   Iter2 80K 86.6 82.5 83.4 71.7 42.3 50.9 69.6↑3.7
RefineCoder-QW-7B   Iter3 80K 87.2 83.3 85.6 72.9 42.0 51.1 70.4↑4.5
Table 2: Results on HumanEval, MBPP and BigCodeBench. Values in red indicate severe data leakage.
Dataset Size HumanEval (+) MBPP (+) LCB BCB
Magic-OSS-Instruct Wei et al. (2024b) 75K 4.45 9.40 2.63 4.56
Magic-Evol-Instruct Wei et al. (2024b) 110K 43.20 19.40 2.91 4.10
Evol-CodeAlpaca-v1 Luo et al. (2024) 110K 47.04 19.46 2.91 4.69
Code-FeedBack Zheng et al. (2024a) 68K 30.50 17.67 3.16 4.49
Ours 80K 4.97 7.00 1.54 5.00
Table 3: The Test Leakage Indicator (TLI) quantifies data leakage by measuring the average maximum n-gram overlap between dataset samples and benchmark samples. The larger the TLI value, the more severe the data leakage.
4.4 Implement Details
Iterative Training We employ DSCoder-6.7B-Base and Qwen2.5-Coder-7B-Base as the base pre-trained models. To obtain an initial code model $M_{0}$, we fine-tune DSCoder-6.7B-Base and Qwen2.5-Coder-Base on the 80K initial SFT dataset, resulting in RefineCoder-DS-6.7B and RefineCoder-QW-7B. The two RefineCoder models undergo three iterative training under our ACR framework. In the ACR pipeline, the model self-samples 7 code responses with a temperature of 0.7. We use Qwen2.5-32B-Instruct Hui et al. (2024) as the Judge and Critic. We set the number of training epochs to 2, the global batch size to 64, and the learning rate to 5e-6, employing the AdamW optimizer with a cosine learning rate decay strategy. All our training uses 16 A100-80G GPUs, utilizing the LLaMA-Factory Zheng et al. (2024b).
Evaluation We use the Pass@1 metric to evaluate the performance of model on benchmarks. We prioritize baseline results from the benchmark leaderboards or the original papers. If unavailable, we evaluate them locally using the same settings as the RefineCoder series.
4.5 Experimental Results
As shown in Table 1, the RefineCoder series achieves impressive performance gains on two challenging code benchmarks. While the initial RefineCoder-DS-6.7B and RefineCoder-QW-7B exhibit moderate performance, the two models achieve an average Pass@1 improvement of 2.7 and 3.1, respectively, after three refinement iterations. Notably, the performance improved by 2.0 and 2.3 points from iter0 to iter1, highlighting that incorporating self-generated code and external critique yields better results than directly distilling code responses from the teacher model. Compared to baselines of the same size, the RefineCoder-DS-6.7B and RefineCoder-QW-7B (Iter3) outperform them by 3.6 and 2.3 points, respectively, using only 80K data. This demonstrates the superiority of our method in improving advanced coding ability.
Similarity, our method also improves the basic coding ability, as shown in Table 2. After three iterations, our model achieves average performance improvements of 2.7 and 4.5 points, respectively, on HumanEval(+), MBPP(+) and BigCodeBench. This demonstrates that the model is capable of identifying and addressing deficiencies by learning from both its self-generated code and external feedback. Compared with all baselines, our models achieve the best average Pass@1 performance. Nevertheless, we find that the performance of RefineCoders on the Humaneval(+) are inferior to the state-of-the-art baseline. This discrepancy prompts us to conduct a data leakage analysis of the datasets in the next section.
5 Further Study
5.1 Data Leakage Analysis
In Table 2, we observe that while the RefineCoder (Iter3) surpasses the baselines on average pass@1, it still lags behind the state-of-the-art baseline on HumanEval(+) and MBPP+. To better explore this performance gap, we investigate potential data leakage in the datasets used by the baselines.
Concretely, we analyze the four open-source datasets  [^1] https://huggingface.co/datasets/ise-uiuc/Magicoder-OSS-Instruct-75K  [^2] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K  [^3] https://huggingface.co/datasets/theblackcat102/evol-codealpaca-v1  [^4] https://huggingface.co/datasets/m-a-p/Code-Feedback with the help of the TLI metric proposed by Wang et al. (2024b). These datasets are used by MagiCoder-S-DS-6.7B, OpenCodeInterpreter-DS-6.7B, and WaveCoder-Ultra-6.7B (The datasets for DSCoder-6.7B-Instruct and Qwen2.5-Coder-7B-Instruct are not publicly available). The results are shown in Table 3, Magic-OSS-Instruct and our dataset are entirely free from data leakage. In contrast, Magic-Evol-Instruct, Evol-CodeAlpaca-v1, and Code-FeedBack demonstrate severe data leakage concerning HumanEval (+) and MBPP (+). These three datasets are used by three baselines, making it challenging for RefineCoder to surpass them. We also present an intuitive scatter plot of similarity scores in the appendix D for a more detailed analysis.
HumanEval+ (%) Best Worst Error Rate
RefineCoder-DS-6.7B Iter3 76.2 60.4-15.8 4.3
RefineCoder-QW-7B Iter3 83.6 75.0-8.6 3.1
MBPP+ (%) Best Worst Error Rate
RefineCoder-DS-6.7B Iter3 70.3 60.1-10.2 5.8
RefineCoder-QW-7B Iter3 73.4 71.1-2.3 5.3
Table 4: Best and Worst denote the performance of code with the highest and lowest scores. Error rate denotes the proportion of cases where the best code fails the test cases but the worst code passes.
5.2 The effectiveness of Composite Scoring System
The performance improvement through iterative training preliminarily validates the effectiveness of our scoring system. In this section, we further verify it by enabling the model to sample 10 code responses for each programming question in benchmarks and scoring them. As shown in Table 4, the performance of the highest-scoring code far exceeds that of the lowest-scoring code, while the scoring system maintains a low error rate. This clearly validates its effectiveness.
Figure 3: Ablation Study. The y-axis denotes the average pass@1 value on LiveCodeBench and BigCodeBench-hard.
Figure 4: Evaluation with feedback using RefineCoder-QW-7B.
5.3 The effectiveness of Selective Critiquing Strategy
We conduct ablation experiments to validate the effectiveness of the selective critiquing strategy in ACR. As shown in Figure 3, for each iteration, we construct two variants of RefineCoder (without 3.1 module of Figure 2 or without 3.2 module). During the iteration process, ablating either module leads to a performance drop in both RefineCoder-QW-7B and RefineCoder-DS-6.7B, indicating the effectiveness of our strategy. Furthermore, removing the second-turn critique data has a greater impact on performance. We believe this action prevents the model from reflecting on self-generated low-quality code and external critique, thereby hindering performance improvement.
5.4 Multilingual Code Generation
Models C++ C# Java Bash TypeScript JavaScript AVG
RefineCoder-DS-6.7B   Iter0 56.5 58.9 58.9 40.5 62.3 64.6 57.2
RefineCoder-DS-6.7B   Iter1 62.1 61.4 55.1 35.4 65.4 66.5 58.2
RefineCoder-DS-6.7B   Iter2 63.4 56.2 56.2 38.6 65.4 65.8 57.5
RefineCoder-DS-6.7B   Iter3 63.4 58.2 57.0 39.2 62.9 65.8 57.8
RefineCoder-QW-7B   Iter0 60.7 72.4 68.0 49.9 77.6 78.1 67.8
RefineCoder-QW-7B   Iter1 62.1 74.6 65.0 50.8 76.8 77.4 67.8
RefineCoder-QW-7B   Iter2 63.7 74.1 69.6 49.7 78.3 77.7 68.9
RefineCoder-QW-7B   Iter3 60.9 75.3 73.8 52.2 79.3 77.9 69.9
Table 5: Performance of RefineCoder on the MultiPL-E.
As shown in Table 5, we also evaluate the out-of-distribution (OOD) code generation performance on the multilingual benchmark MultiPL-E Cassano et al. (2022) despite fine-tuning on a Python-only dataset. After one and three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B achieve optimal average performance, surpassing the initial model by 1.0 and 2.1 points, respectively. In particular, RefineCoder-QW-7B (Iter3) achieves the best results among half of the programming languages. This demonstrates that the effectiveness of our ACR method generalizes well to the OOD code generation domain.
5.5 Evaluation with External Feedback
Iterative ACR not only improves the one-off code generation performance of RefineCoder but also endows it with the ability to correct errors based on feedback. Following  Zheng et al. (2024a), we design two types of external feedback to evaluate this ability of RefineCoder: 1) Execution Feedback: Model refines its self-generated erroneous code based on execution results from the executor; 2) Human-like Feedback: GPT-4o first analyzes the programming question, initial error code, and execution results to generate improvement suggestions that mimic human thinking. The model then refines the code based on these suggestions. The results of RefineCoder-QW-7B on HumanEval(+) and MBPP(+) are shown in Figure 4. It can be observed that external feedback improves performance on all benchmarks, with human-like feedback yielding a more significant enhancement. Surprisingly, with the help of human-like improvement suggestions, RefineCoder-QW-7B (Iter1 and Iter3) outperforms GPT-4o on three benchmarks, demonstrating its ability to understand feedback and refine code.
6 Conclusions
In this paper, we propose the Adaptive Critique Refinement (ACR) method to iteratively refine code LLMs using self-generated code and external critique, which breaks away from the traditional teacher distillation paradigm and improves the intrinsic one-off code generation ability. We design a composite scoring system and a selective critiquing strategy. These two components are centered around LLM-as-a-Judge and LLM-as-a-Critic to evaluate and critique self-generated code. This simulates the process where a student solves a problem independently and then refines it by comparing it with the reference answer. We develop the RefineCoder-DS-6.7B and RefineCoder-QW-7B models and demonstrate the effectiveness of iterative ACR on multiple code generation benchmarks. Further studies reveal the impact of the each component in ACR.
7 Limitations
Although Adaptive Critique Refinement (ACR) and the associated RefineCoder demonstrate significant improvements in the code generation task, several limitations remain. First, ACR is primarily designed to refine code responses based on programming instructions, necessitating an initial set of high-quality and diverse instructions. Therefore, a specialized code instruction generator is still required to make ACR more automated. Furthermore, while ACR can apply to other reasoning-intensive tasks, such as mathematics, this paper has not fully explored these domains.
Appendix A Prompts for Judge and Critic
The prompt for point/pair-wise LLM-as-a-Judge and LLM-as-a-Critic are shown in Figure 8,  9 and  10.
Appendix B Prompts for Constructing SFT Dataset
We called GPT-4o to create the SFT dataset, with the following prompts used:
$\circ$ Prompt for generating code-related concepts and programming instructions, as shown in Figure 11.
$\circ$ Prompts for evolving existing instructions, as shown in Figures 12,  13,  1415161718,  19,  20.
$\circ$ For generating code responses, We directly feed the instructions to the model.
Appendix C Prompts for Evaluation
For Humaneval (+) and MBPP (+), we use the prompts designed by OpenCodeInterpreter Zheng et al. (2024a); for Livecodebench and Bigcodebench-hard, we use the official prompts. When conducting evaluations with execution feedback, we have designed a prompt as illustrated in Figure 21. When executing human-like feedback, following OpenCodeInterpreter, we generate two prompts: one for generating improvement suggestions and the other for refine code according to human-like feedback, as shown in Figures 22 and  23.
Appendix D Data Leakage Analysis
We use 4-grams and 5-grams to compute the similarity between the instructions in the training set and the questions in the benchmark. And then we draw scatter plots by selecting the top-200 data points with the highest similarity from each dataset, as shown in Figure 7. The scatter plot visually illustrates the data leakage phenomenon in the dataset, primarily concentrated in the HumanEval (+) and the latter half of the MBPP (+). This makes it difficult for our model to surpass baselines on these two benchmarks. In contrast, on the leakage-free LiveCodeBench and BigCodeBench-hard benchmarks, our model can outperform all baselines.
Appendix E Diminishing Returns
As shown in Figure 5, we present the proportions of two types of data during the iterative process of RefineCoder-QW-7B. The self-generated code struggles more to outperform as iterations continue, indicating diminishing potential for further gains through self-refinement. The corresponding experimental results are shown in Figure 6. As observed, the model exhibits clear diminishing returns after the third iteration. This suggests that the model may have entered a mild overfitting stage, where further refinement not only fails to improve performance but may also exacerbate the issue.
Figure 5: Proportions of two types of data in the iterative process.
Figure 6: The average Pass@1 value of RefineCoder-QW-7B on HumanEval(+), MBPP(+) and BigCodeBench.
Figure 7: Scatter plot of similarity scores between the datasets and four benchmarks, with average similarity scores for different datasets indicated by horizontal dashed lines. We selected the top 200 data points with the highest similarity from each dataset. Different y-axis ranges are set for better visualization.
Figure 8: The prompt template for pointwise judgment.
Figure 9: The prompt template for pairwise judgment.
Figure 10: The prompt template for pairwise critique.
Figure 11: The prompt template for generating code-related concepts and instructions.
Figure 12: The prompt template for Addition evolution.
Figure 13: The prompt template for Breath evolution.
Figure 14: The prompt template for Complexity evolution.
Figure 15: The prompt template for Concretizing evolution.
Figure 16: The prompt template for Deepening evolution.
Figure 17: The prompt template for Diversion evolution.
Figure 18: The prompt template for Increase evolution.
Figure 19: The prompt template for Misdirection evolution.
Figure 20: The prompt template for Reason evolution.
Figure 21: The prompt template for Execution Feedback.
Figure 22: The prompt template for generating improvement suggestions.
Figure 23: The prompt template for Human-like Feedback.